OpenStudy (anonymous):

Sean wants to make a mixture that is 50% lemon juice and 50% lime juice. How much 100% lemon juice should he add to a juice that is 30% lemon juice and 70% lime juice to make 7 gallons of the 50% lemon/50% lime juice mixture?

OpenStudy (anonymous):

A. 0.5 gallon B. 1 gallon C. 2 gallons D. 2.5 gallons

OpenStudy (nurali):

p= # of pure lemon juice m= # of gallons of 30/70 mixture 7= total # of gallons of 50/50 mix p+m=7 Amount of lemon juice in p gallons of 100% lemon juice = 1p Amount of lemon juice in m gallons of 30% lemon juice = .3m Amount of lemon juice in 7 gallons of 50% lemon juice = 3.5 so p+.3m=3.5 Now solve: p+m=7 p=7-m substitute 7-m into second equation for p (7-m) +.3m=3.5 -.7m=-3.5 m=5 substitute 5 for m into the equation p=7-m p=7-5 p=2 gallons of 100% lemon juice for wanted results